A cat walks from the neighbors house to your house which is 15.0 m in 45 sec. What is the cat's speed?

Select one:

a. 675 m/s

b. 3.00 m/s

c. 0.33 m/s

d. 0.944 m/s

A ball with a horizontal speed of 1.5 m/s rolls off a bench that is 2.2 m high. How long will it take to reach the floor and how far will the ball land away from the bench?

Select one or more:

a. .45 sec

b. 1.47 sec

c. 2.15 sec

d. .67 sec
Answer #1

here,

a)

time taken by cat , t = 45 s

distance , d = 15 m

the cat's speed , v = d/t

v = 15/45 = 0.33 m/s

the cat's speed is 3.33 m/s

b)

horizontal speed of the ball , vx = 1.5 m/s

height , h = 2.2 m

let the time taken to reach the ground be t

h = uy * t + 0.5 * g * t^2

2.2 = 0 + 4.9 * t^2

t = 0.67 s

the time taken is d) 0.67 s

the horizontal distance , x = vx * t = 1 m